A refrigerator working on reversed Carnot cycle requires 0.5 KW 

per KW of cooling to maintain a temperature of -15°C. Determine the 

following:

a) COP of the refrigerator

b) Temperature at which heat is rejected and

c) Amount of heat rejected to the surroundings per KW of cooling.


1
Expert's answer
2021-06-14T14:09:20-0400

Answer:-

"W=0.5kw\\
Q_1=1.0kw\\
T_2=-15^o C + 273 =258K\\"


a) COP = "\frac{1.0}{0.5}=2"

b) so, cop = "\frac{T_2}{T_1-T_2}"

"2=\frac{258}{T_1-258}\\
T_1=387K=114^oC"


c)Amount of heat rejected to srrounding

"Q_1=W_1+Q_2=0.5+1
\\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ =1.5 KW"